Grade upto college level Physical Chemistry

A 1.0 g sample of Fe2O3 solid of 55.2% purity is dissolved in acid and reduced by heating the solution with zinc dust. The resultant solution is cooled and made upto 100.0 ml. An aliquot of 25.0 ml of this solution requires 17.0 ml of 0.0167 M solution of an oxidant for titration. Calculate the number of electrons taken up by the oxidant in the reaction of the above titration.

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
Mass of Fe2O3 in the sample = 55.2/100 * 1 = 0.552 g
Number of moles of Fe2O3 = 0.552/159.8 = 3.454 * 10-3
Number of moles of Fe3+ ions = 2 * 3.454 * 10-3
= 6.9 * 10-3mol = 6.90 mmol
Since its only 1 electron is exchanged in the conversion of Fe3+ to Fe2+, the molecular mass is the same as equivalent mass.
∴ Amount of Fe2+ ion in 100 ml. of sol. = 6.90 meq
Volume of oxidant used for 100 ml of Fe2+ sol
= 17 * 4 = 68 ml.
Amount of oxidation used = 68 * 0.0167 mmol
= 1.1356 mmol
Let the number of electrons taken by the oxidant = n
∴ No. of meq.of oxidant used = 1.1356 * n
Thus 1.1356 * n = 6.90
N = 6.90/1.1356 = 6